# UNO-Bench: A Unified Benchmark for Exploring the Compositional Law Between Uni-modal and Omni-modal in OmniModels

### v2

@@ 3.3.2 General Scoring Model
Beside the dataset construction, multi-step open-ended question introduces a new challenge of automated evaluation. To overcome this obstacle, we propose a general scoring model that supports multi-choice question, single-step open-ended question and multi-step open-ended question at the same time. Since the task is to compare the target answer and the predicted answer, we use Qwen3-14B(Yang et al., 2025) as backbone and curate a training dataset as illustrated in Figure.5. One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types. See details in Appendix.D. Through the human-in-the-loop dataset curation, the scoring model achieves 95% accuracy in out-of-distribution models and benchmarks.
@@
Figure 4: Example of multi-step open-ended questions.
Figure 5: Training dataset construction and model training pipeline for general scoring model.
Table 2: Omni-modal benchmark results on multimodal models.
@@
Table 3: Omni-modal analysis on model ability taxonomy. Cross-modal Recognition refers to the set of other Perception capabilities except Cross-modal Alignment.
@@
4 Experiment and Analysis
4.1 Experiment Setting
We evaluate omni models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-Omni-30B-A3B-Instruct(Xu et al., 2025b), Qwen-2.5-Omni-3B, Qwen-2.5-Omni-7B(Xu et al., 2025a), Baichuan-Omni-1.5(Li et al., 2025), MiniCPM-O-2.6(Yao et al., 2024), and Ming-lite-Omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025). To have a fair comparison between instruct model and thinking model, we adopt similar way in Qwen-3(Xu et al., 2025b) that limits thinking budget to 128 tokens. We apply this restriction to Gemini-2.5-Pro and disable the thinking mode for both Gemini-2.5-Flash and Gemini-2.0-Flash. All the other model integrations strictly adhere to official implementations. In video processing, each model receives raw video and performs frame sampling according to its own sampling strategy.
In the following sections, we conduct comprehensive experiments on UNO-Bench and try to answer the following questions:
  1. 1.
    How about the performance of current omni models and their shortage abilities?
  2. 2.
    What is the relationship of uni-modal ability and omni-modal ability?
  3. 3.
    Could the UNO-Bench provide an effective evaluation for the omni model?
4.2 Model Performance
4.2.1 Overall Analysis
@@ 4.2.2 Top-tier Analysis
Why Gemini-2.5-Pro performances significantly better than other models? We attempt to provide an analysis and some hypotheses. On one hand, it stems from the leading uni-modal understanding ability. On the other hand, regarding the reasoning mechanism, Gemini is equipped with audio captioning functionalities as indicated in the technical report(Comanici et al., 2025), and illustrated in Figure.6. It can also naturally incorporate audio content as a foundation for reasoning. Existing open-source models, due to their smaller size, lack reasoning processes in a multimodal context. Limited reasoning mostly relies on text or images, with few involving specific audio content.
@@
However, does Gemini performance good enough? To answer this question, we invited human experts to conduct a competition. It is important to note that, unlike the dataset annotators, the invited human experts have not been exposed to the question or the answer.
Finding 1. Gemini-2.5-Pro has reached human comparable perception ability in omni-modal understanding, yet there remains a gap in its reasoning performance. Compared to human experts, Gemini-2.5-Pro only has 8.3% less performance, which shows humanity comparable intelligence. The comparison of scores for specific ability items can be seen in Figure.7. Upon examining ability analysis, we observe an intriguing phenomenon: humans are more proficient in reasoning as opposed to perception (81.3% compared to 74.3%), which contrasts with the model’s performance. By interviewing various human experts, it becomes evident that humans might miss some information in video or audio formats, and their world knowledge is more limited compared to large language models.
@@
We arrive at the following result through rigorous mathematical derivation, and the detailed derivation process is provided in the Appendix B.
@@
We then posit a fusion symmetry assumption: in end-to-end omni models, the fusion mechanism does not inherently favor one modality over anotherXu et al. (2025b); Yao et al. (2024), implying symmetric scaling behavior. This leads to $\alpha=\beta$. Substituting this into Eq. 10, we arrive at the Omni-modal Compositional Law:
@@
The exponent $\alpha\approx 2.19$ is the most critical discovery, revealing a powerful Power-law synergy. Because $\alpha>1$, the function is convex, meaning the rate of performance gain accelerates. This explains the transition from a "short-board effect" to an "emergent ability" seen in Figure.8:
  • Limited Gains at Low Performance: For models with weaker uni-modal abilities (e.g., MiniCPM-O), the curve is relatively flat. Small improvements in the product of uni-modal scores yield only marginal gains in omni-modal performance. This can be seen as a "short-board effect", where the system is not yet capable of effectively leveraging the combined inputs.
@@
Table 4: Ablation of visual understanding ability.
@@ 4.3.2 Ablation Audio Understanding
To isolate the impact of auditory information, we evaluated models under three conditions: visual-only (Visual), visual plus transcribed audio (+Caption), and the full audio-visual input (+Audio). We further divided the audio into two categories: Environment sounds, for which we provided textual descriptions, and Speech, for which we provided ASR transcripts. All transcriptions were generated by a powerful multimodal model. The results are presented in Table.5.
@@
Table 5: Ablation of audio understanding ability.
@@ 4.4.2 Dataset Compression
To verify the compression performance, we define these metrics: Spearman’s Rank Correlation Coefficient(SRCC) for ranking consistency, Pearson’s Linear Correlation Coefficient(PLCC) for linear value consistency, Root Mean Square Error(RMSE) for numerical precision, Margin of Error(MoE) for quantifying estimation uncertainty, and Confidence Interval Coverage(CIC) for statistical reliability.
@@ 4.4.3 Benchmark Comparison
To ensure the quality of dataset, we conduct quality check on 10%-20% random samples in each benchmarks. As shown in Table.1, UNO-Bench has 100% accuracy on omni-modal dataset while 98% questions requires cross-modality to solve. It shows the highest quality among existing omni model benchmarks.
@@ 5 Conclusion
In this work, we introduce a high quality and diversity benchmark to evaluate omni models comprehensively. The main foundings are summarized as below.
1. With unified data framework in UNO-Bench, we found that the multimodal capability may not simply be a linear superposition of uni-modalities, but rather follows a significant multiplicative relationship. The evaluation results show that it manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
2. Both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni models. The Gemini-2.5-Pro shows comparable perception capability with human experts but still has a performance gap in reasoning aspect.
3. Besides better dataset quality and evaluation efficiency, UNO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
In the future work, we will extend the dataset’s scale by human-in-the-loop automated pipeline and hold a private test set to avoid hacking. The ability coverage also needs to extend to more difficult reasoning tasks like STEM and code. At the same time, the relationship among cross-modals understandings and how to improve them are still exciting problem to explore. Furthermore, our compositional law has only been validated on UNO-Bench with uniformly distributed capabilities. Whether this law still holds under different capability distributions remains to be explored.
@@ Appendix A Benchmarks Utilized in Dataset Compression
  • General visual question answering, RealWorldQAxAI (2023), MMEChaoyou et al. (2023), SeedBenchLi et al. (2023).
  • Document and chart understanding, OCRBench Liu et al. (2024d), FoxLiu et al. (2024a), DocLocal4kYe et al. (2023).
  • Stem & reasoning, MMMUYue et al. (2024), MMMU-ProYue et al. (2025), CMMMUGe et al. (2024), MathVistaLu et al. (2024b), MathVisionWang et al. (2024b), ScienceVistaTeam et al. (2025), GMAI-MMBenchYe et al. (2024).
  • Multi-image Understanding, ReMiKazemi et al. (2024), MuirBenchWang et al. (2024a).
  • Audio question answering, MMAUSakshi et al. (2025), MMSUWang et al. (2025), SDQAFaisal et al. (2021).
@@ Appendix C Model Selection for the Compositional Law
Appendix D General Scoring Model
One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types. We define six different question types in Figure.12.
Appendix E Model Ability Taxonomy
@@
Appendix F Omni-modal Solubility and Uni-modal Solubility Cases
Omni-modal solubility means that both visual and audio information are required to answer the question. While uni-modal solubility means that only visual or audio information is required to answer the question.
There are two uni-modal solubility cases. One is when audio and video present the same amount of information, and only one modality is needed to solve the problem, as shown in Figure.17(b). The other is when one modality provides useful information, while the other modality is unhelpful, as shown in Figure.17(c). The omni-modal solubility case can be seen in Figure.17(a).
Figure 17: Comparison of the problem solvable by each modality.

### v3

@@ 3.3.2 General Scoring Model
Beside the dataset construction, multi-step open-ended question introduces a new challenge of automated evaluation. To overcome this obstacle, we propose a general scoring model that supports multi-choice question, single-step open-ended question and multi-step open-ended question at the same time. Since the task is to compare the target answer and the predicted answer, we use Qwen3-14B(Yang et al., 2025) as backbone and curate a training dataset as illustrated in Figure.6. One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types, as shown in Figure.7. Through the human-in-the-loop dataset curation, the scoring model achieves 95% accuracy in out-of-distribution models and benchmarks.
@@
Figure 5: Construction of Multi-Step Open-Ended Questions.
Figure 6: Training pipeline for general scoring model.
4 Experiment and Analysis
4.1 Experiment Setting
We evaluate omni models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-Omni-30B-A3B-Instruct(Xu et al., 2025a), Qwen-2.5-Omni-3B, Qwen-2.5-Omni-7B(Xu et al., 2025b), Baichuan-Omni-1.5(Li et al., 2025), MiniCPM-O-2.6(Yao et al., 2024), and Ming-lite-Omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025). To have a fair comparison between instruct model and thinking model, we adopt similar way in Qwen-3(Xu et al., 2025a) that limits thinking budget to 128 tokens. We apply this restriction to Gemini-2.5-Pro and disable the thinking mode for both Gemini-2.5-Flash and Gemini-2.0-Flash. All the other model integrations strictly adhere to official implementations. In video processing, each model receives raw video and performs frame sampling according to its own sampling strategy.
In the subsequent sections, we perform detailed experiments on UNO-Bench and aim to address the following questions:
  1. 1.
    How do current omni models perform, and what are their limitations?
  2. 2.
    How are uni-modal and omni-modal capabilities related?
  3. 3.
    Is the UNO-Bench capable of effectively evaluating the omni model?
4.2 Model Performance
4.2.1 Overall Analysis
Table 2: General performance of omni models in UNO-Bench for both uni-modal capability and omni-modal capability, where omni-modal benchmark includes multi-choice questions (Omni-MC) and multi-step open-ended questions (Omni-MO).
@@ 4.2.1 Overall Analysis
Table 3: Analysis of Omni-MC on ability taxonomy. To simplify the analysis, Cross-modal Recognition refers to the set of other Perception capabilities except Cross-modal Alignment.
@@ 4.2.2 Top-tier Analysis
What makes the performance of Gemini-2.5-Pro stand out compared to other models? We aim to offer an analysis along with several hypotheses. On one hand, it stems from the leading uni-modal understanding ability. On the other hand, regarding the reasoning mechanism, Gemini is equipped with audio captioning functionalities as indicated in the technical report(Comanici et al., 2025), and illustrated in Figure.9. It can also naturally incorporate audio content as a foundation for reasoning. Existing open-source models, due to their smaller size, lack reasoning processes in a multimodal context. Limited reasoning mostly relies on text or images, with few involving specific audio content.
@@
The successive question is whether Gemini’s performance measures up. To answer this question, we invited human experts for a competition. It’s crucial to highlight that, unlike the dataset annotators, these experts had no prior exposure to the questions or answers.
Finding 1. Gemini-2.5-Pro has reached human comparable perception ability in omni-modal perception, yet there remains a gap in its reasoning performance. Compared to human experts, Gemini-2.5-Pro exhibits similar performance in perception, but falls behind in reasoning. The comparison of scores for specific ability items can be seen in Figure.9. Upon examining ability analysis, we observe an intriguing phenomenon: humans are more proficient in reasoning as opposed to perception (81.3% compared to 74.3%), which contrasts with the model’s performance. By interviewing various human experts, it becomes evident that humans might miss some information in video or audio formats, and their world knowledge is more limited compared to large language models.
@@
We arrive at the following result through rigorous mathematical derivation, and the detailed derivation process is provided in the Appendix.B.
@@
We then posit a fusion symmetry assumption: in end-to-end omni models, the fusion mechanism does not inherently favor one modality over another (Xu et al., 2025a; Yao et al., 2024), implying symmetric scaling behavior. This leads to $\alpha=\beta$. Substituting this into Eq. 2, we arrive at the Omni-modal Compositional Law:
@@
The exponent $\alpha\approx 2.19$ is the most critical discovery, revealing a powerful Power-law synergy. Because $\alpha>1$, the function is convex, meaning the rate of performance gain accelerates. This explains the transition from a ”short-board effect” to an ”emergent ability” seen in Figure.10:
  • Limited Gains at Low Performance: For models with weaker uni-modal abilities (e.g., MiniCPM-O), the curve is relatively flat. Small improvements in the product of uni-modal scores yield only marginal gains in omni-modal performance. This can be seen as a ”short-board effect”, where the system is not yet capable of effectively leveraging the combined inputs.
@@
4.3.1 Ablation Visual Understanding
Table 4: Ablation of visual understanding ability. The three settings are audio-only (Audio), audio plus high-quality textual captions of the visual scene (+Caption), and the full audio-visual input (+Visual).
@@ 4.3.2 Ablation Audio Understanding
To isolate the impact of auditory information, we evaluated models under three conditions: visual-only (Visual), visual plus transcribed audio (+Caption), and the full audio-visual input (+Audio). We further divided the audio into three categories: the Speech category was annotated with ASR transcripts, while both the Environment and Music categories received textual descriptions. To ensure the robustness of our analysis and improve statistical reliability, the data-insufficient Music class was merged with the Environment class. The majority of the transcriptions were manually produced by human annotators, while a smaller subset was generated by a powerful multimodal model. The results are presented in Table.5.
@@
Table 5: Ablation of audio understanding ability. The three settings are visual-only (Visual), visual plus transcribed audio (+Caption), and the full audio-visual input (+Audio). We further divided the audio into two categories: Environment sounds, for which we provided textual descriptions, and Speech, for which we provided ASR transcripts.
@@ 4.4.3 Benchmark Comparison
To ensure the quality of dataset, we conduct quality check on 10%-20% random samples in each benchmarks. As shown in Table.1, UNO-Bench has 100% accuracy on omni-modal dataset while 98% questions requires cross-modality to solve. It shows the highest quality among existing omni benchmarks.
@@ 5 Conclusion
In this work, we introduce a high quality and diversity benchmark to evaluate omni models comprehensively. With unified data framework in UNO-Bench, we found that the omni-modal capability may not simply be a linear superposition of uni-modal capabilities, but rather follows a significant multiplicative relationship. The evaluation results show that it manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models. In addition, we found that both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni models. The Gemini-2.5-Pro shows comparable perception capability with human experts but still has a performance gap in reasoning aspect. Besides better dataset quality and evaluation efficiency, UNO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
In the future work, we will extend the dataset’s scale by the human-in-the-loop automated pipeline and hold a private test set to avoid hacking. The ability coverage also needs to extend to more difficult reasoning tasks like STEM and code. At the same time, the relationship among cross-modals understandings and how to improve them are still exciting problem to explore. Furthermore, our compositional law has been validated on UNO-Bench with uniformly distributed capabilities. Whether this law still holds under different capability distributions remains to be explored.
@@ Appendix A Benchmarks Utilized in Dataset Compression
  • General visual question answering, RealWorldQAxAI [2023], MMEChaoyou et al. [2023], SeedBenchLi et al. [2023].
  • Document and chart understanding, OCRBench Liu et al. [2024b], FoxLiu et al. [2024d], DocLocal4kYe et al. [2023].
  • Stem & reasoning, MMMUYue et al. [2024], MMMU-ProYue et al. [2025], CMMMUGe et al. [2024], MathVistaLu et al. [2024b], MathVisionWang et al. [2024a], ScienceVistaTeam et al. [2025], GMAI-MMBenchYe et al. [2024].
  • Multi-image Understanding, ReMiKazemi et al. [2024], MuirBenchWang et al. [2024c].
  • Audio question answering, MMAUSakshi et al. [2025], MMSUWang et al. [2025], SDQAFaisal et al. [2021].
@@ Appendix C Model Selection for the Compositional Law
Appendix D Model Ability Taxonomy